Circular No. 8971

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 R2 (SCOTTI) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery

of comet P/2001 X2 (cf. IAUCs 7775, 7777) on Spacewatch CCD images obtained with the 1.8-m f/2.7 reflector at Kitt Peak, the object showing a coma diameter of 5′′ and a tail extending 0′.5 in p.a. 272^{o}. The indicated correction to the prediction on MPC 56802 is ΔT = −0.15 day.

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.5.43015  4^{h}35^{m}05.^{s}42  +20^{o}39′24′′.5  20.5

Incidental astrometry from the Mt. Lemmon survey on 2007 Sept. 13 (mag 20.8–22.3) and 2007 Oct. 8 (mag 21.6–21.9), and from the Spacewatch survey on 2008 Aug. 7, are included with the current astrometry on MPEC 2008-R30, together with revised orbital elements and an ephemeris.

V466 ANDROMEDAE H. Yamaoka, Kyushu University, reported the discovery by K. Itagaki

(Yamagata, Japan) of a possible nova on a CCD image taken on Sept. 1.6 UT with a 0.21-m reflector in the course of his sky survey, with a confirming unfiltered CCD image taken on Sept. 1.603 with a 0.60-m reflector that yields mag about 12.7 and the following precise position: α = 2^{h}00^{m}25^{s}.40, δ = +44^{o}10′18′′.7 (equinox 2000.0). A patrol image from Jan. 26.532 shows nothing brighter than mag 17.5 at this position; there is a very faint star (red mag 20.4, blue mag 21.2) in the GSC 2.3 catalogue with position end figures 25^{s}.44, 19′′.1. Additional confirming observations were published on CBET 1491, including investigations that found nothing brighter than blue mag ∼ 20.5 at the position of this variable. S. Korotkiy, Moscow, and colleagues report that photometry shows a light curve with amplitude 0.13 mag and period 0.055(1) day. N. N. Samus, Institute of Astronomy, Moscow, informs us that this variable has been assigned the designation V466 And.

P. Challis and R. P. Kirshner, Harvard-Smithsonian Center for Astrophysics; and P. Garnavich, University of Notre Dame, report that a spectrum (range 350-730 nm) of V466 And was obtained with the 1.5-m Tillinghast telescope (+ FAST spectrograph) on Sept. 3.497 UT, showing a strong blue continuum with Hα and Hβ in emission. Later Balmer lines are a combination of broad absorption and weak emission. The FWHM of Hα is 600 km/s. Also seen are He I 492.1-, 587.5-, 667.8-, 501.5-, and 706.5-nm and He II 486.1-nm. The spectrum is that of a dwarf nova in outburst.

2008 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green